# Demonstrating dynamic surface codes

### v1

@@ Abstract
Quantum Error Correction (QEC) enables accurate quantum computation in the presence of uncontrolled physical noise [9]. The surface code [1, 10, 11] is a preeminent example, capable of suppressing physical errors below a relatively high threshold and providing excellent capabilities for logical computation [2, 12]. These desirable characteristics have inspired extensive experimental work on the surface code [13, 14], including recently demonstrated below-threshold performance on a superconducting quantum processor [15] and transversal logical gates in a neutral atom quantum processor [16]. To date, all of these experiments have been designed around a square grid of statically assigned qubits implementing CNOT/CZ entangling interactions to match the standard circuit implementation. This was assumed to comprise the minimum requirement to perform the code, since fault-tolerant circuits are traditionally constructed from static QEC codes.
Recent progress in the theory of time-dynamic QEC, including dynamic codes [4] and space-time stabilizers [17, 18], challenge the traditional static stabilizer approach and unlock new opportunities for the construction of fault-tolerant circuits. The new framework of the space-time detecting region generalizes these approaches, defined by local and connected regions of space-time that report the presence of errors. This framework permits the re-interpretation of QEC circuits as intermeshed detecting regions, featuring overlapping sensitivity such that all physical errors can be appropriately triangulated. It also highlights that maintaining the precise shape of such detecting regions is less crucial than ensuring all circuit locations are appropriately covered. Exploiting this freedom opens the door to an expansive equivalence class of circuits for implementing well-understood QEC codes [5, 19, 6, 20, 7], providing distinct flexibility in their hardware implementation.
@@
The newfound freedom in circuit implementation can be understood naturally using detecting regions, which indicate all locations in a QEC circuit where a non-commuting error would be detected by measurements the region touches. Fig. 1a shows a standard distance-5 surface code patch, where we highlight a single detecting region in Fig. 1c. This Z-type detecting region is sensitive to any X or Y error occurring in highlighted locations in the circuit. The region propagates through the circuit in the same way that Pauli terms propagate through Clifford operations, illustrated in Fig. 1b. The detecting region persists over two consecutive cycles, first expanding from a single qubit to cover the static code stabilizer immediately after the first measurement gate, which we call the end-cycle state, then contracting to end in a second measurement. Time-slices of the detecting region are shown in Fig. 1d; using a tiling of spacetime through overlapping detecting regions, we ensure that all circuit locations are checked for errors. Given this general framework, circuit flexibility arises by modifying the gate layers and associated detecting regions while preserving the tiling of spacetime.
@@
Using a time-dynamic circuit featuring an alternative detecting region tiling, a recent proposal provides an embedding of the surface code on a hexagonal grid [5]. In Fig. 2a we highlight the key insights of this proposal. The first insight is that the last two-qubit gate layer of each QEC cycle can be altered to match first layer with the CNOT orientation inverted, using the same qubit-qubit coupling layer twice in each cycle. This results in the expanding detecting region being shifted laterally while preserving the square weight-4 pattern of stabilizers. The second key insight is to use a time-reversed circuit every second cycle, refocusing the detecting region to its starting position during the contraction by an opposite lateral shift. This pattern of two opposing detecting regions is sufficient to tile the space-time bulk of the circuit. We include details on the construction of detecting regions at the spatial boundaries in (see SM sec. C1).
@@
We run the memory experiment for a distance-5 hexagonal surface code and five tiling distance-3 codes. In Fig. 2b we report the logical error probability at varying circuit depths, along with the logical error rate per cycle for different decoders (Fig. 2c), ordered from fastest to most accurate. Using the Harmony decoder [29] with a noise model prior trained using reinforcement learning [30], we measure a distance-5 logical error rate of and an mean distance-3 logical error rate of averaged over five codes. From this, we extract an error suppression factor of , indicating the surface code can be operated on a hexagonal lattice with performance matching the standard square lattice implementation () [15].
@@
Walking is permitted by the gate layers and detecting region time-slices shown in Fig. 3a. Using this walking circuit, we benchmark the scaling of a logical memory on a 105-qubit superconducting quantum processor [15], with logical error probability results shown in Fig. 3b and logical error rate per cycle for different decoders in Fig. 3c. Here, we realize the distance-5 and embedded distance-3 walking circuits without any explicit leakage removal strategy (no DQLR). In order to allow room for walking, the distance-5 circuit uses 58 qubits, more than the standard circuit’s 49, leading to a slight overhead in logical error rate. For a logical memory experiment, we measure distance-5 and average distanced-3 logical error rates of , , leading to an error suppression factor . The median CZ Pauli error of used in the walking circuit is larger than the hexagonal implementation (see SM sec. D2), leading to a reduction in .
@@
To benchmark the walking circuit’s leakage dynamics, we measure the leakage population at each time-slice of the circuit, separated into the two sub-grids of qubits as shown in Fig. 3d. At each moment when a reset is applied to a group of qubits, the leakage population on that group drops by around an order in magnitude, followed by a ramp in population until they are reset again two cycles later. The staggered reset applications limit leakage population buildup over many cycles, preventing a rising contribution to the overall logical performance as the code continues.
@@
Leakage also causes time-correlated errors that are not properly handled by the decoder. We directly probe these correlations by measuring the detector autocorrelation in Fig. 3e, which is the average covariance of the detecting events separated by n cycles. While two-round correlations are expected from time-like Pauli errors (e.g. readout or reset errors), this measurement reveals longer time correlations induced by non-Pauli errors. Without leakage removal in the standard code, we measure a long correlation tail lasting nearly 50 cycles, while the walking circuit and the standard code with leakage removal remove these long-time correlations.
@@
In addition to these leakage measurements, we apply the same detector budgeting techniques introduced in the previous section to reveal the contributions of each physical error mechanism on detection probabilities in the circuit (Fig. 3f). Our Pauli model produces detections that have a weaker correlation with experiment (RMS , see (see SM sec. D3)) than the hexagonal implementation, indicating that some error mechanisms are not fully captured by the Pauli model, but reveals the same dominant contributions as the hexagonal implementation. The CZ contribution is larger as expected from the reduced gate performance. Even with the slight reduction in logical error rate and , our results demonstrate that a walking surface code can suppress errors as distance increases with a built-in leakage removal, paving the way for new architectures that can use this property to their advantage.
@@
In Fig. 4b and Fig. 4c, we benchmark the logical error rate of a distance-5 and embedded distance-3 iSWAP implementation of surface codes on our 72-qubit device (the distance-5 circuit uses 57 qubits). Using decoder H*, we measure logical error rates of and averaged over the five distance 3 codes, here again without using DQLR. We extract the error suppression factor , demonstrating the viability of a surface code implemented with iSWAP gates.
@@
We probe the effect of leakage during memory experiments in Fig. 4h by measuring the population of excited states following the injection of a state on a single bulk data qubit as shown in Fig. 4g. We repeat this experiment for three surface code implementations: the standard CZ implementation with and without DQLR, and the iSWAP implementation (which does not use DQLR). In a reference experiment where no leakage is injected (dotted lines), the standard surface code without DQLR reaches a steady state with oscillating average leakage between and per qubit due to the large leakage generation of the CZ gates. With DQLR, the average leaked population is considerably reduced as expected, oscillting between and . With the iSWAP implementation, the average leakage population reaches an oscillating steady state between and , limiting leakage well below the standard implementation without DQLR, indicating the superior leakage properties of the iSWAP gate compared with the CZ gate. Of all three, the standard CZ code with DQLR leads to the best leakage performance.
@@
To compare surface code implementations on current hardware, in Fig. 5a we show contributions of each error component to the inverse scaling factor, , which can be expanded as a linear sum in physical error rates at first order (see SM sec. B). These simulations go beyond the Pauli models used to construct the detector error budgets by including effects such as leakage and crosstalk. For all codes, these leakage and crosstalk contributions are still small relative to the two qubit gates, dynamical decoupling (DD), and readout error. The DD contribution is largest for the iSWAP code, indicating the more complex layout of data and measure qubits end-cycle moments (see SM sec. E1) can add additional error.
For the hexagonal implementation, while the performance we measure already matches state of the art, tailoring the device to a honeycomb or brickwall lattice will reduce the number of connections needed from 2 to 1.5 per qubit (taking into account that qubits share connections). In principle, it will also reduce the number of potential error channels by reducing the complexity of the device. From an frequency optimization perspective, the lower connectivity reduces the number of variables, simplifying the optimization step. In Fig. 5b, we simulate the potential gain of tailoring the hardware to a hexagonal lattice by accounting for realistic device inhomogeneities in hardware parameters as well as TLS density with a frequency optimization step [27] (see SM sec. G). From these simulations we predict that simplifying the hardware by removing one coupler per qubit can lead to a improvement of the scaling factor . Such an increase in compounds favorably with a reduction in number of couplers needed, simplifying the design and number of control lines.
@@ Acknowledgements
Alec Eickbusch, Matt McEwen, Volodymyr Sivak1, Alexandre Bourassa1, Juan Atalaya1, Jahan Claes1, Dvir Kafri1, Craig Gidney1, Christopher W. Warren1, Jonathan Gross1, Alex Opremcak1, Nicholas Zobrist1 Kevin C. Miao1, Gabrielle Roberts1, Kevin J. Satzinger1, Andreas Bengtsson1, Matthew Neeley1, William P. Livingston1, Alex Greene1, Rajeev Acharya1, Laleh Aghababaie Beni1, Georg Aigeldinger1, Ross Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Abraham Asfaw1, Ryan Babbush1, Brian Ballard1, Joseph C. Bardin1, 6, Alexander Bilmes1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Michael Broughton1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Josh Cogan1, Roberto Collins1, Paul Conner1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Sayan Das1, Alexander Del Toro Barba1, Sean Demura1, Laura De Lorenzo1, Agustin Di Paolo1, Paul Donohoe1, Ilya K. Drozdov1, 4, Andrew Dunsworth1, Aviv Moshe Elbag1, Mahmoud Elzouka1, Catherine Erickson1, Vinicius S. Ferreira1, Leslie Flores Burgos1, Ebrahim Forati1, Austin G. Fowler1, Brooks Foxen1, Suhas Ganjam1, Gonzalo Garcia1, Robert Gasca1, Élie Genois1, William Giang1, Dar Gilboa1, Raja Gosula1, Alejandro Grajales Dau1, Dietrich Graumann1, Tan Ha1, Steve Habegger1, Monica Hansen1, Matthew P. Harrigan1, Sean D. Harrington1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Jeremy Hilton1, Hsin-Yuan Huang1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Zhang Jiang1, Xiaoxuan Jin1, Cody Jones1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Hui Kang1, Amir H. Karamlou1, Kostyantyn Kechedzhi1, Trupti Khaire1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Bryce Kobrin1, Alexander N. Korotkov1, Fedor Kostritsa1, John Mark Kreikebaum1, Vladislav D. Kurilovich1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Kim-Ming Lau1, Justin Ledford1, Kenny Lee1, Brian J. Lester1, Loïck Le Guevel1, Wing Yan Li1, Alexander T. Lill1, Aditya Locharla1, Erik Lucero1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Cameron Maxfield1, Jarrod R. McClean1, Seneca Meeks1, Anthony Megrant1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Ramis Movassagh1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alex Pizzuto1, Rebecca Potter1, Orion Pritchard1, Chris Quintana1, Ganesh Ramachandran1, Matthew J. Reagor1, David M. Rhodes1, Eliott Rosenberg1, Elizabeth Rossi1, Kannan Sankaragomathi1, Henry F. Schurkus1, Michael J. Shearn1, Aaron Shorter1, Noah Shutty1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alex Sztein1, Douglas Thor1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Grayson Young1, Adam Zalcman1, Yaxing Zhang1, Ningfeng Zhu1, Sergio Boixo1, Julian Kelly1, Vadim Smelyanskiy1, Hartmut Neven1, Dave Bacon1, Zijun Chen1, Paul V. Klimov1, Pedram Roushan1, Charles Neill1, Yu Chen1, Alexis Morvan.

### v2

@@ Abstract
Quantum Error Correction (QEC) enables accurate quantum computation in the presence of uncontrolled physical noise [9]. The surface code [1, 10, 11] is a preeminent example, capable of suppressing physical errors below a relatively high threshold and providing excellent capabilities for logical computation [2, 12]. These desirable characteristics have inspired extensive experimental work on the surface code [13, 14], including recently demonstrated below-threshold performance on a superconducting quantum processor [15] and transversal logical gates in a neutral atom quantum processor [16]. To date, all of these experiments have been designed around a square grid of statically assigned qubits implementing CNOT/CZ entangling interactions to match the standard circuit implementation. This was assumed to comprise the least demanding experimental implementation of the code, since fault-tolerant circuits are traditionally constructed from static QEC codes.
Recent progress in the theory of time-dynamic QEC, including dynamic codes [4] and space-time stabilizers [17, 18], challenge the traditional static stabilizer approach and unlock new opportunities for the construction of fault-tolerant circuits. The new framework of the space-time detecting region generalizes these approaches, defined by local and connected regions of space-time that report the presence of errors. This framework permits the re-interpretation of QEC circuits as intermeshed detecting regions, featuring overlapping sensitivity such that all physical errors can be appropriately triangulated. It also highlights that maintaining the precise shape of such detecting regions is less crucial than ensuring all circuit locations are covered. Exploiting this freedom opens the door to an expansive equivalence class of circuits for implementing well-understood QEC codes [5, 19, 6, 20, 7], providing distinct flexibility in their hardware implementation.
@@
The newfound freedom in circuit implementation can be understood naturally using detecting regions, which indicate all locations in a QEC circuit where a non-commuting error would be detected by measurements the region touches. Fig. 1a shows a standard distance-5 surface code patch, where we highlight a single detecting region in Fig. 1c. This Z-type detecting region is sensitive to any X or Y error occurring in highlighted locations in the circuit. The region propagates through the circuit in the same way that Pauli terms propagate through Clifford operations, illustrated in Fig. 1b, with more details given in [5]. The detecting region persists over two consecutive cycles, first expanding from a single qubit to cover the static code stabilizer immediately after the first measurement gate, which we call the end-cycle state, then contracting to end in a second measurement. Time-slices of the detecting region are shown in Fig. 1d; using a tiling of spacetime through overlapping detecting regions, we ensure that all relevant circuit locations are checked for errors. Given this general framework, circuit flexibility arises by modifying the gate layers and associated detecting regions while preserving the tiling of spacetime.
@@
Using a time-dynamic circuit featuring an alternative detecting region tiling, a recent proposal provides an embedding of the surface code on a hexagonal grid [5]. In Fig. 2a we highlight the key insights of this proposal. The first insight is that the last two-qubit gate layer of each QEC cycle can be altered to match the first layer with the CNOT orientation inverted, using the same qubit-qubit coupling layer twice in each cycle. This results in the expanding detecting region being shifted laterally while preserving the square weight-4 pattern of stabilizers. The second key insight is to use a time-reversed circuit every second cycle, refocusing the detecting region to its starting position during the contraction by an opposite lateral shift. This pattern of two opposing detecting regions is sufficient to tile the space-time bulk of the circuit. We include details on the construction of detecting regions at the spatial boundaries in (see SM sec. C1).
@@
We run the memory experiment for a distance-5 hexagonal surface code and five tiling distance-3 codes. In Fig. 2b we report the logical error probability at varying circuit depths, along with the logical error rate per cycle for different decoders (Fig. 2c), ordered from fastest to most accurate. Using the Harmony decoder [29] with a noise model prior trained using reinforcement learning [30], we measure a distance-5 logical error rate of and a mean distance-3 logical error rate of (averaged over five codes), fit with the same procedure described in [15]. From this, we extract an error suppression factor of , indicating the surface code can be operated on a hexagonal lattice with performance matching the standard square lattice implementation () [15].
@@
Walking is permitted by the gate layers and detecting region time-slices shown in Fig. 3a. Using this walking circuit, we benchmark the scaling of a logical memory on a 105-qubit superconducting quantum processor [15] (a different device than was used for the hexagonal implementation), with logical error probability results shown in Fig. 3b and logical error rate per cycle for different decoders in Fig. 3c. Here, we realize the distance-5 and embedded distance-3 walking circuits without any explicit leakage removal strategy other than multi-level reset (no DQLR). In order to allow room for walking, the distance-5 circuit uses 58 qubits, more than the standard circuit’s 49, leading to a slight overhead in logical error rate. For a logical memory experiment, we measure distance-5 and average distanced-3 logical error rates of , , leading to an error suppression factor . The median CZ Pauli error of used in the walking circuit is larger than the hexagonal implementation (see SM sec. D2), leading to a reduction in .
@@
To benchmark the walking circuit’s leakage dynamics, we measure the leakage population at each time-slice of the circuit, separated into the two sub-grids of qubits as shown in Fig. 3d. At each moment when a reset is applied to a group of qubits, the leakage population on that group drops by around an order in magnitude, followed by a ramp in population until they are reset again two cycles later. The staggered multi-level reset applications limit leakage population buildup over many cycles, preventing a rising contribution to the overall logical performance as the code continues.
@@
Leakage also causes time-correlated errors that are not properly handled by the decoder. We directly probe these correlations by measuring the detector autocorrelation in Fig. 3e, which is the average covariance of the detecting events separated by cycles. While two-round correlations are expected from time-like Pauli errors (e.g. readout or reset errors), this measurement reveals longer time correlations induced by non-Pauli errors. Without leakage removal in the standard code, we measure a long correlation tail lasting nearly 50 cycles, while the walking circuit and the standard code with leakage removal significantly reduce these long-time correlations.
@@
In addition to these leakage measurements, we apply the same detector budgeting techniques introduced in the previous section to reveal the contributions of each physical error mechanism on detection probabilities in the circuit (Fig. 3f). Our Pauli model produces detections that have a weaker correlation with experiment (RMS , see (see SM sec. D3)) than the hexagonal implementation, indicating that some error mechanisms are not fully captured by the Pauli model, but reveals the same dominant contributions as the hexagonal implementation. The CZ contribution is larger as expected from the reduced gate performance. Even with the reduction in logical performance and , our results demonstrate that a walking surface code can suppress errors as distance increases with a built-in leakage removal, paving the way for new architectures that can use this property to their advantage.
@@
In Fig. 4b and Fig. 4c, we benchmark the logical error rate of a distance-5 and embedded distance-3 iSWAP implementation of surface codes on our 72-qubit device (the distance-5 circuit uses 57 qubits). Using decoder H*, we measure logical error rates of and averaged over the five distance 3 codes, here again without using DQLR (relying on multi-level reset as the only active source of leakage removal). We extract the error suppression factor , demonstrating the viability of a surface code implemented with iSWAP gates.
@@
We probe the effect of leakage during memory experiments in Fig. 4h by measuring the population of excited states following the injection of a state on the center-most bulk data qubit after 10 cycles as shown in Fig. 4g, here using the distance-5 code. We repeat this experiment for three surface code implementations: the standard CZ implementation with and without DQLR, and the iSWAP implementation (which does not use DQLR). In a reference experiment where no leakage is injected (dotted lines), the standard surface code without DQLR reaches a steady state with oscillating average leakage between and per qubit due to the large leakage generation of the CZ gates. With DQLR, the average leaked population is considerably reduced as expected, oscillting between and . With the iSWAP implementation, the average leakage population reaches an oscillating steady state between and , limiting leakage well below the standard implementation without DQLR, indicating the superior leakage properties of the iSWAP gate compared with the CZ gate. Of all three, the standard CZ code with DQLR leads to the best leakage performance.
@@
To compare surface code implementations on current hardware, in Fig. 5a we show contributions of each error component to the inverse scaling factor, , which can be expanded as a linear sum in physical error rates at first order (see SM sec. B). The total cycle times of the implementations are similar (hex: , walking: , iSWAP: ), allowing for direct comparison. These simulations go beyond the Pauli models used to construct the detector error budgets by including effects such as leakage and crosstalk. For all codes, these leakage and crosstalk contributions are still small relative to the two qubit gates, dynamical decoupling (DD), and readout error. The cphase contribution for the iSWAP code is significant, accounting for of the budget, and causes the iSWAP code’s simulated to be larger than the other three implementations.
For the hexagonal implementation, while the performance we measure already matches state of the art, tailoring the device to a honeycomb or brickwall lattice will reduce the number of connections needed from 2 to 1.5 per qubit (taking into account that qubits share connections). In principle, it will also reduce the number of potential error channels by reducing the complexity of the device. From a frequency optimization perspective, the lower connectivity reduces the number of variables, simplifying the optimization step. In Fig. 5b, we simulate the potential gain of tailoring the hardware to a hexagonal lattice by accounting for realistic device inhomogeneities in hardware parameters as well as TLS density with a frequency optimization step [27] (see SM sec. G). From these simulations we predict that simplifying the hardware by removing one coupler per qubit can lead to a improvement of the scaling factor . Such an increase in compounds favorably with a reduction in number of couplers needed, simplifying the design and number of control lines.
@@ Acknowledgements
Alec Eickbusch, Matt McEwen, Volodymyr Sivak1, Alexandre Bourassa1, Juan Atalaya1, Jahan Claes1, Dvir Kafri1, Craig Gidney1, Christopher W. Warren1, Jonathan Gross1, Alex Opremcak1, Nicholas Zobrist1, Kevin C. Miao1, Gabrielle Roberts1, Kevin J. Satzinger1, Andreas Bengtsson1, Matthew Neeley1, William P. Livingston1, Alex Greene1, Rajeev Acharya1, Laleh Aghababaie Beni1, Georg Aigeldinger1, Ross Alcaraz1, Trond I. Andersen1, Markus Ansmann1, Frank Arute1, Kunal Arya1, Abraham Asfaw1, Ryan Babbush1, Brian Ballard1, Joseph C. Bardin1, 3, Alexander Bilmes1, Jenna Bovaird1, Dylan Bowers1, Leon Brill1, Michael Broughton1, David A. Browne1, Brett Buchea1, Bob B. Buckley1, Tim Burger1, Brian Burkett1, Nicholas Bushnell1, Anthony Cabrera1, Juan Campero1, Hung-Shen Chang1, Ben Chiaro1, Liang-Ying Chih1, Agnetta Y. Cleland1, Josh Cogan1, Roberto Collins1, Paul Conner1, William Courtney1, Alexander L. Crook1, Ben Curtin1, Sayan Das1, Alexander Del Toro Barba1, Sean Demura1, Laura De Lorenzo1, Agustin Di Paolo1, Paul Donohoe1, Ilya K. Drozdov1, 2, Andrew Dunsworth1, Aviv Moshe Elbag1, Mahmoud Elzouka1, Catherine Erickson1, Vinicius S. Ferreira1, Leslie Flores Burgos1, Ebrahim Forati1, Austin G. Fowler1, Brooks Foxen1, Suhas Ganjam1, Gonzalo Garcia1, Robert Gasca1, Élie Genois1, William Giang1, Dar Gilboa1, Raja Gosula1, Alejandro Grajales Dau1, Dietrich Graumann1, Tan Ha1, Steve Habegger1, Michael C. Hamilton1, 4, Monica Hansen1, Matthew P. Harrigan1, Sean D. Harrington1, Stephen Heslin1, Paula Heu1, Oscar Higgott1, Reno Hiltermann1, Jeremy Hilton1, Hsin-Yuan Huang1, Ashley Huff1, William J. Huggins1, Evan Jeffrey1, Zhang Jiang1, Xiaoxuan Jin1, Cody Jones1, Chaitali Joshi1, Pavol Juhas1, Andreas Kabel1, Hui Kang1, Amir H. Karamlou1, Kostyantyn Kechedzhi1, Trupti Khaire1, Tanuj Khattar1, Mostafa Khezri1, Seon Kim1, Bryce Kobrin1, Alexander N. Korotkov1, Fedor Kostritsa1, John Mark Kreikebaum1, Vladislav D. Kurilovich1, David Landhuis1, Tiano Lange-Dei1, Brandon W. Langley1, Kim-Ming Lau1, Justin Ledford1, Kenny Lee1, Brian J. Lester1, Loïck Le Guevel1, Wing Yan Li1, Alexander T. Lill1, Aditya Locharla1, Erik Lucero1, Daniel Lundahl1, Aaron Lunt1, Sid Madhuk1, Ashley Maloney1, Salvatore Mandrà1, Leigh S. Martin1, Orion Martin1, Cameron Maxfield1, Jarrod R. McClean1, Seneca Meeks1, Anthony Megrant1, Reza Molavi1, Sebastian Molina1, Shirin Montazeri1, Ramis Movassagh1, Michael Newman1, Anthony Nguyen1, Murray Nguyen1, Chia-Hung Ni1, Logan Oas1, Raymond Orosco1, Kristoffer Ottosson1, Alex Pizzuto1, Rebecca Potter1, Orion Pritchard1, Chris Quintana1, Ganesh Ramachandran1, Matthew J. Reagor1, David M. Rhodes1, Eliott Rosenberg1, Elizabeth Rossi1, Kannan Sankaragomathi1, Henry F. Schurkus1, Michael J. Shearn1, Aaron Shorter1, Noah Shutty1, Vladimir Shvarts1, Spencer Small1, W. Clarke Smith1, Sofia Springer1, George Sterling1, Jordan Suchard1, Aaron Szasz1, Alex Sztein1, Douglas Thor1, Eifu Tomita1, Alfredo Torres1, M. Mert Torunbalci1, Abeer Vaishnav1, Justin Vargas1, Sergey Vdovichev1, Guifre Vidal1, Catherine Vollgraff Heidweiller1, Steven Waltman1, Jonathan Waltz1, Shannon X. Wang1, Brayden Ware1, Travis Weidel1, Theodore White1, Kristi Wong1, Bryan W. K. Woo1, Maddy Woodson1, Cheng Xing1, Z. Jamie Yao1, Ping Yeh1, Bicheng Ying1, Juhwan Yoo1, Noureldin Yosri1, Grayson Young1, Adam Zalcman1, Yaxing Zhang1, Ningfeng Zhu1, Sergio Boixo1, Julian Kelly1, Vadim Smelyanskiy1, Hartmut Neven1, Dave Bacon1, Zijun Chen1, Paul V. Klimov1, Pedram Roushan1, Charles Neill1, Yu Chen1, Alexis Morvan.
@@
Department of Electrical and Computer Engineering, Auburn University, Auburn, AL, USA